### 16 Feb 2024 v2

# Machine Learning, Density Functional Theory, and Experiments to Understand the Photocatalytic Reduction of CO by CuPt/TiO

Abstract
The photoconversion of CO to hydrocarbons is a sustainable route to its transformation into value-added compounds and, thereby, crucial to mitigating the energy and climate crises. CuPt nanoparticles on TiO surfaces have been reported to show promising photoconversion efficiency. For further progress, a mechanistic understanding of the catalytic properties of these CuPt/TiO systems is vital. Here, we employ ab-initio calculations, machine learning, and photocatalysis experiments to explore their configurational space and examine their reactivity and find that the interface plays a key role in stabilizing *CO, *CO, and other CH-containing intermediates, facilitating higher activity and selectivity for methane. A bias-corrected machine-learning interatomic potential trained on density functional theory data enables efficient exploration of the potential energy surfaces of numerous CO@CuPt/TiO configurations via basin-hopping Monte Carlo simulations, greatly accelerating the study of these photocatalyst systems. Our simulations show that CO preferentially adsorbs at the interface, with C atom bonded to a Pt site and one O atom occupying an O-vacancy site. The interface also promotes the formation of *CH and *CH intermediates. For confirmation, we synthesize CuPt/TiO samples with a variety of compositions and analyze their morphologies and compositions using scanning electron microscopy and energy-dispersive X-ray spectroscopy, and measure their photocatalytic activity. Our computational and experimental findings qualitatively agree and highlight the importance of interface design for selective conversion of CO to hydrocarbons.
keywords:
Machine Learning Interatomic Potential, Basin-Hopping Monte Carlo, Density Functional Theory, Titania, Metal/Oxide Interface, Photocatalysis, CO Activation, CO Conversion1 Introduction
CO capture and its efficient conversion into value-added products are important steps to mitigating the energy and climate crises. CO-derived chemicals like polycarbonates, urea etc. and fuels like methane, alcohols, jet fuel etc. have industrial applications, and thus can provide routes for monetization. Owing to its natural abundance, low-operating costs, high-chemical stability, low toxicity, and environmental compatibility, TiO and its derived materials have been extensively investigated over roughly seven decades, with several promising applications [1, 2, 3, 4] including catalysis [5, 6, 7, 8, 9] and photocatalysis.[10, 11, 12, 13] Using mechanistic insights to rationally optimize TiO-based materials is fundamental to enhance CO-capture efficiency and improve activity and selectivity for its conversion to hydrocarbons.
Despite the challenges of low-yield ( mol) for large-scale applications in photocatalytic CO conversion to hydrocarbons, the activity and selectivity of TiO-based systems [14, 15, 13, 16] are encouraging. Since pristine TiO suffers from a low yield due to several reasons, including the maximum-visible light under-utilization,[17] large electron-hole separation,[10] and wide band gaps 3.0 eV (rutile)[18] and 3.2 eV (anatase),[19] several strategies have been tested insofar. Manipulating the geometric/electronic properties of TiO via creation of oxygen vacancies or depositing metal (e.g. Cu/Pt) nanoparticles have potential to enhance the photocatalytic activity.[20]
On TiO surfaces, CO activation is a bottleneck step owing to a variety of reasons.[21, 22, 23, 24] For instance, the *CO formation via absorption of photo-excited electrons: CO *CO+e is thermodynamically hindered whilst the *CHOO formation via absorption of two electrons: CO+H *CHOO+2e is kinetically hindered.[24] The trivial charge transfer between *CO and TiO would result in weaker interaction.[21, 22, 23] Since CO weakly interacts with defect-laden TiO (see section S1 and Fig.S1 in supporting information) it cannot facilitate CO activation without external driving force, thus necessitating an engineering of TiO structures.
The manipulation of geometrical and electronic structures of TiO-based systems has remarkable impacts on their chemical activity. Under reducing conditions, a TiO surface contains O vacancies which bolster metal-support interaction and enhance its reactivity.[25, 26] The Cu-decorated TiO promotes CO conversion.[27] The suitable band realignment due to charge transfer,[28] enhances activity of Pt/TiO for CO conversion to CH.[29] The PtCu/TiO shows better activity than Pt/TiO.[30] CuPt/titania offers a good activity towards the photo-catalytic CO conversion and higher selectivity for CH (yield 92%).[13] Due to a low percentage, the TiO(110) might be mostly exposed as a reactive facet. In this work, we model CuPt/TiO systems by considering the rutile TiO(110) with O vacancy and 13-atom CuPt clusters.
For the accelerated discovery of materials for CO reduction, various machine learning (ML) models have been utilized to study several systems [31, 32, 33, 34] including (bare) semiconductor oxides.[32] More recently, E(3)-equivariant neural network (ENN) is attracted attention owing to high data efficiency for ML models.[35, 36] The ML approach in combination with basin-hopping Monte Carlo (BHMC) method can provide an efficient way to tackle the combinatorial problem of composition of bimetallic Cu-Pt nanoclusters together with CO adsorbates, relating to the large degrees of freedom due to several adsorption sites and the orientation of clusters and CO.
Whilst the ab-initio methods offer atomistic insights into the potential energy surface (PES) of material systems but suffer from the high computational cost and inefficient scaling with system sizes, ML interatomic potentials (MLIPs) can provide substantial speed-ups at the expense of accuracy with linear scaling. ML can often reduce computational costs by orders of magnitude.[37, 38] MLIPs have been improved drastically to achieve higher accuracy in energies ( meV/atom) and forces (0.1 eV/Å) and have been utilized in several applications, e.g. in understanding complex materials chemistry,[39, 40, 41, 42, 43, 44], enumerating nanostructures configurations,[41, 45] performing long-time MD simulations,[46, 47, 48] and exploring catalytic reactions.[49, 50, 51, 52] Nevertheless, MLIPs suffer from highly inhomogeneous feature-space sampling of training set (inherent biases). The inclusion of underrepresented configurations leads to significant errors since the data representing inhomogeneous local environments are overshadowed by core atoms.[53] To model CO@CuPt/TiO systems, it thus requires an inherent-bias correction.
Herein, we perform the ab-initio calculations of CuPt/TiO(110) model systems to generate the atom coordinates, energy, and forces for the initial training of MLIP based on ENN.[36] We modify the algorithm to reduce unwanted errors in under-represented configurations by using the Gaussian density function-based weighting scheme, improving MLIP reliability and transferability. We use the unbiased MLIP together with BHMC algorithm to enumerate several configurations of CO@CuPt/TiO and find that CO adsorbs at interfacial sites. CO activation is then validated by DFT with calculations of adsorption energy, CO bond length/angle changes, and charge transfer. Having examined the key reaction pathways for overall reaction between CO and HO, we provide mechanistic insights into the role of CuPt/TiO interface in CO reduction activity. Our simulations results qualitatively agree with photocatalysis experiments.
2 Results
2.1 MLIP training and validation
In Fig.1, we show the parity plot of MLIP-predicted against DFT-calculated energy and forces, along with the distribution of errors (inset). For the training set (8,646 structures), the root-mean-squared error (RMSE) and mean-absolute error (MAE) in energy are 0.63 and 0.46 meV/atom, respectively, while those in forces are 0.06 and 0.04 eV/Å, respectively. The model is also tested using a validation set (9,633), resulting in the equivalent RMSE and MAE errors.
To obtain a better understanding of MLIP errors to learn ab-initio data, we obtain decomposed-force parities for CO, CuPt cluster, and TiO (Fig.S2). Even with less representation in the overall training data, force prediction errors for both CO and nanocluster are comparable to the overall RMSE of force distribution (Fig.S2). The effect of weighted training has been discussed further in the supporting information (Fig.S3).
2.2 Potential energy surface exploration for CO adsorption on CuPt/TiO
We perform BHMC+MLIP simulations to efficiently determine low-energy configurations of CO@CuPt/TiO. Analyzing the relative energy of CO@CuPt/TiO, e.g. (Fig.2), the BHMC-explored PES curve reveals that initially explored structures (b-c) are higher in energy which evolve towards minima region structures (e-f). If lower-energy structures are not found, the algorithm gradually accepts higher-energy configurations, moving to higher-energy basins (g-h).
The exploration yields the lowest energy basins with a common feature i.e. CO adsorption occurs at CuPt/TiO interface. We perform similar simulations with different Cu/Pt cluster compositions and discuss the minima CO adsorption structures (section 2.4). We provide details of exploration procedure in SI (see section-S2.2). Despite MLIP’s low validation errors, comparing structural stability within a small energy range is not viable. We adopt a 0.1 eV energy window, adding BHMC-generated structures with energy differences less than 0.1 eV to the low-energy ensemble. Subsequently, DFT relaxation is applied to this ensemble to eliminate the potential small MLIP errors.
The extremely large-configurational space is apparent from an illustration of PES exploration for a model system (Fig.2), whose configurations are found to be different cluster shapes with various arrangement of Pt/Cu atoms together with several possible adsorption sites for CO, including the CuPt/TiO interfacial sites as well as different oxygen vacancy sites on TiO(110).
2.3 Interaction of CuPt nanocluster with TiO
For CuPt composition, the cluster-support interaction is facilitated mostly by chemical bonds formed between Cu and bridging O atoms and bonds between Pt and bridging O and/or five-fold () coordinated Ti atoms. The distributions of Cu-O bond lengths for the lowest-energy configurations (Fig.S4) and more details can be found in SI (see S2.3). Owing to the metal-support interaction, the CuPt cluster adopts a hemi-spherical shape (Fig.3). Regardless of the composition, the supported CuPt cluster prefers a 3D geometrical shape.
It is important to understand the distribution of CuPt/TiO frontier states since the states spatially distributed around interfacial atoms would play an important role in ultrafast dynamics and thereby photo-induced reactions.[54] For CuPt/TiO (), the frontier states are spattially localized at interface and around Cu-Pt atoms (Fig.3). Although the spatial localization is found to vary with Cu-Pt composition, the common feature is that there is some distribution of maximum valence band (VB) frontier states. In contrast, Cu/TiO VB frontier states localize at the TiO second layer (from the top). These results suggest that the CuPt/TiO interface would be more active than Cu/TiO interface.
2.4 CO interaction with CuPt/TiO
We now examine the interaction of CO with CuPt/TiO via DFT by considering the lowest-minima configurations predicted by MLIP+BHMC simulations. For all CuPt compositions, we find that CO adsorption is thermodynamically favorable, and CO is activated (Table 1). Regardless of the CuPt composition, CO adsorbs at the interface and the degree of interaction depends upon the composition and shape of clusters. The CO interaction with Pt/TiO is stronger as compared to Cu/TiO systems. Although we do not observe any specific trend in adsorption energies, our calculations suggest that the CO interaction magnitude would be higher for CuPt clusters than for only pure Cu. As compared to the bond length (1.162 Å) and angle (180) of CO in gas phase, upon its adsorption the C-O bond elongates, and the bond angle changes to 121-134. Taken together results of the CO interaction with CuPt/TiO systems with various Pt/Cu compositions (=-0.50 to -1.2 eV), the charge transfer to an adsorbed CO (0.7 to 0.9) and the elongation of C-O bond (0.2 Å) and the substantial reduction of OCO (55-59), we show that CuPt/TiO systems facilitate CO activation.
| Pt | Adsorption | d | OCO | d | Charge gained | |
| site | d | by CO | ||||
| 0 | Cu O | -0.42 | 1.266, 1.321 | 121.4 | 1.940 | 1.00 |
| 1 | Cu O | -0.62 | 1.212, 1.370 | 122.9 | 1.921 | 0.87 |
| 2 | Pt O | -1.24 | 1.271, 1.306 | 121.3 | 2.006 | 0.95 |
| 3 | Pt O | -0.71 | 1.218, 1.320 | 127.7 | 2.028 | 0.78 |
| 4 | Pt O | -0.67 | 1.217, 1.315 | 128.3 | 2.020 | 0.75 |
| 5 | Pt O | -0.95 | 1.253, 1.287 | 125.5 | 1.997 | 0.79 |
| 6 | Pt O | -0.80 | 1.216, 1.352 | 124.1 | 2.021 | 0.80 |
| 7 | Pt Ti | -0.86 | 1.246, 1.251 | 134.6 | 1.977 | 0.54 |
| 8 | Pt O | -0.94 | 1.216, 1.364 | 123.0 | 2.020 | 0.82 |
| 9 | Pt O | -1.15 | 1.232, 1.358 | 122.5 | 1.987 | 0.96 |
| 10 | Pt Ti | -0.48 | 1.309, 1.217 | 128.2 | 2.016 | 0.64 |
| 11 | Pt O | -0.67 | 1.350, 1.215 | 124.6 | 2.017 | 0.79 |
| 12 | Pt Ti | -1.03 | 1.240, 1.265 | 133.6 | 1.961 | 0.53 |
| 13 | Pt Ti | -1.01 | 1.242, 1.266 | 134.4 | 1.951 | 0.58 |
[…]
Fig.4 and others in Fig.S5 (SI, S2.4) present the lowest-minima structures of CO@CuPt/TiO configurations. CO adsorbs at the interface of Cu/TiO such that C attaches to a Cu atom and O atom fills out the O vacancy site (Fig.4a). For CuPt with (Pt 1), the CO adsorbs at the interface with an O atom occupying O vacancy site and C bonded to Pt, thus preferring a Pt over Cu site (Fig.4b-e). For , CO adsorbs in a way that C is bonded to Pt and a O to Ti (Fig.4f). The CO adsorption geometries in the case of are like that of . These results indicate that the CuPt/TiO interface serves as active sites for CO adsorption, particularly highlighting the importance of interfacial Cu/Pt and surface Ti atoms (near O vacancy).
From the electron density difference () plot (Fig.4g), it is seen that the resulting electron density is re-distributed in the bonding regions of *CO atoms (C, O1, O2) and interfacial atoms (Pt/Ti). Thus, the favorable interaction between CO and CuPt/TiO is mediated by the interface. The charge re-distribution, upon CO interaction with CuPt/TiO, suggests that the charge transfer takes place between two sub-systems. There is charge accumulation around O atoms and charge depletion around a C atom and surface Ti atoms. Some charge addition to the *CO anti-bonding state can trigger the C-O bond elongation. Indeed, based on the partial charges analysis, the net charge transfer of 0.96e takes place from CuPt/TiO to an adsorbed CO (Table 1), inducing the C-O bond elongation and thereby facilitating CO activation.
Hereinafter, we present the results for structure and energetics properties of CuPt/TiO model system when interacting with the reactants (CO/HO) and key intermediates of CO reduction. The model selection is partly based on the stronger CO adsorption at interfacial sites of CuPt/TiO and partly on the higher Pt concentration than Cu (prevent a CuPt cluster from oxidizing).
2.5 Energetics for CO dissociation vs hydrogenation
Based on the DFT-calculated energetics, the *CO formation from *CO dissociation is energetically favorable (E=-0.62 eV). In contrast, the CO hydrogenation via *CO+*H *CHOO+* step is endothermic (E=+0.60 eV). Similarly, the *CO+*H*COOH+* is endothermic (E=+0.27 eV). Although these steps require some energy, the reactions might be activated in normal reaction conditions. Calculated energetics provide some indication that while *CO formation is a favorable process (*CO hydrogenation can proceed to yield *CHO formation, being more thermodynamically favorable than *COH formation.[55]), the *CO hydrogenation might also feasible under external driving force.
2.6 Structures and energetics of intermediates
In Fig.5, we provide the energy profile for reaction between CO and HO (reactants) and some key intermediates adsorbed on CuPt/TiO. The relative energy for all adsorbed species (with/without co-adsorption) is given in reference to the sum of energies of gas-phase CO and HO, and of CuPt/TiO. The reaction energy (E) is the difference in energy between initial and final states. The adsorption of HO in the presence of *CO is favorable (E=-0.79 eV). The HO adsorbs at a Ti site with H being closer to a bridging O atom. The TiO(110) is well studied for HO activation.[6, 56, 57, 48, 9] The CO adsorption is energetically favorable (E=-1.15 eV). Based on CO activation, we are mainly interested in the chemical activity at/near the interface.
Once *CHOO is formed via hydrogenation step, then it can undergo self-dissociation to *CHO and *O (Fig.5) with the gain of energy (E=-0.21eV). As an alternative pathway, the *CO dissociation can occur with *CO formation. The *CO dissociation (in co-adsorption with *H at a bridging O atom) is more exothermic (E=-0.62 eV) than from *CO dissociation (with *H at a Cu-Pt site) (E=-0.35 eV).
Not only of *CO and *CO, but also the stabilities of other intermediates such as *CH, *CH, *CH are important for the selectivity for CH over CO. Now we examine the structure and energetics of *CHOH (which can be formed via reaction between *H and *CHO, which can be formed either by *CHOO dissociation [27] or by a *H+*CHO step.[27, 58] The CHOH adsorbs at interfacial sites where C attaches to Pt (2.034 Å) and OH functional group occupies an O-vacancy site (Fig.6a). The dissociation step *CHOH *CH+*OH is energetically favorable (E=-0.65 eV) and is better than Pt/TiO (E=-0.12 eV,[27]) thus indicating that *CH formation is thermodynamically feasible and *CH is stabilized by interfacial Cu/Pt atoms. These atoms can serve as active sites for *CH formation and possibly its further reactions. Since the Pt-H distance (H of *OH) is only 2.322 Å, the Pt site can facilitate the reaction between *CH and *OH, to form *CH and *O. For feasibility of this process, however, the *OH bond scission and an additional C-H bond formation should take place simultaneously. The resulting *CH then can react with *H to form *CH or CH (Fig.S7). The *CH+*HCH step can be feasible at room temperature. We note that *CH formation route could also proceed with the reaction between *H and *CH, which is stabilized at the interfacial Cu and Pt atoms, and that *CH formation from the reaction between *H and *CH intermediates. Indeed, the formation of *CH from dissociation of *CHOH, which is also activated at the interface, is energetically much favorable (E=-1.28 eV) (Fig.6b).
2.7 Experimental rationalization of CO photoconversion
We characterize the morphology and composition of reduced P25 samples by scanning electron microscopy (Fig.7a) and energy-dispersive X-ray spectroscopy (Fig.S8), respectively. The reduced P25 is irregular, with a particle size 50 nm. We observe some high-contrast features after photodeposition reaction of Pt and Cu. The Pt detected is similar with input Pt amount. In contrast, the Cu detected is significantly less than the input Cu amount with 0.14Pt-0.6Cu and 1.26Pt-0.4Cu requiring 2.5% and 1.25% Cu, respectively. This incomplete photodeposition of Cu compared to Pt might be due to the lower reduction potential of Cu than Pt and the redox nature of Cu.
We perform photocatalytic reactor measurements under a flow of humidified CO and Ar, illuminated by an AM1.5 filtered Xe lamp focused to 5 sun intensity. In Fig.7b, we show the CH yield over time of photocatalysis reactions on CuPt/P25 samples with various CuPt compositions. The sustained CH production is the highest for a 0.14Pt-0.6Cu sample (with the photodeposition of 0.14 mol% Pt and 2.5 mol% Cu onto the reduced P25), in agreement with an earlier experiment.[13] The low Pt and high Cu composition yields higher CH. The CH yield was higher initially for 1.79Pt-0Cu samples, but over time it became deactivated. Reduced Cu-P25/Pt-P25 or reduced P25 (pure) samples produce significantly less CH, suggesting that the higher CH yield is the consequence of the synergistic effect of Cu, Pt, and P25. The CO photoconversion activity decreases substantially as both Pt and Cu loadings are increased, attributed to the growth of large-sized CuPt particles and the reduction of interfacial density of Cu/Pt-TiO sites, suggesting the TiO-Pt/Cu interface is highly preferable for achieving good photocatalytic conversion of CO to hydrocarbon.
3 Discussion
Compared to oxide supports [21, 23] and metal surfaces,[59, 60] CO is chemically well-activated by metal/oxide interfaces. From our DFT simulations, the defect-laden TiO(110) does not facilitate CO activation ( -0.2 eV), with an onset desorption temperature 175 K.[59] Since TiO itself does not stabilize CO,[59] CO reduction is not feasible. On well-defined metal surfaces, CO adsorption is also weak, e.g. 0.21-0.28 eV for Cu [61] and 0.03 eV for Pt [62] and thus limits to trivial charge transfer between CO and Cu/Pt catalysts. Due to weak interaction with (=-0.29 eV),[62] it can also desorb before its reaction with H.[27] When supported using TiO, metal/oxide systems can enhance CO adsorption. While Pt/TiO(110) binds CO with =-0.61 eV,[27] Cu/TiO(110) does with =-0.65 eV.[63] As compared to bare TiO or Cu/TiO, the CuPt/TiO can be considered as a better catalyst owing to: 1) the higher *CO stability at the interface, and 2) the interfacial Cu and Pt atoms providing sites for nucleophilic and electrophilic adsorbates.
For the catalytic conversion of CO into hydrocarbons, the atomic hydrogens (from HO splitting process) are required to be present on the catalyst surfaces so that they can react directly with either *CO reactant or *CO/*C to form CH-containing intermediates. The CuPt/TiO interface provides active sites for formation of intermediates: *CO, *CHOO, *CHO, *CHOH, *CH, *CH, and *CH, thus highlighting the importance of interfacial sites in the chemical activity of CuPt/TiO systems towards the CO conversion to hydrocarbons.
We propose the reaction pathways as follows: CO adsorbs at the CuPt/TiO interface with CO oxygen at a O vacancy site, and HO adsorbs at a Ti site of the reduced P25, thus facilitating the formation of *H* and *CO at the interfacial Pt sites and leaving behind a surface O. Then, they can react to form intermediates required for CH formation pathway. Finally, the O-vacancies generation is assisted by the redox activity of CuPt nanoparticles with some fraction of Cu on their surface.
Although MLIP/DFT simulations and experimental observations cannot be directly matched, our studies together can provide meaningful insights into CuPt/TiO photocatalytic systems. Their excellent chemical activity can be correlated with the thermodynamically favorable, moderate-to-strong interaction (=-0.67 to -1.24 eV) with CO and the net charge gained by CO (). The synergistic effect of the surface TiO, Cu, and Pt atoms at the interface results in stronger CO adsorption with =-0.67 to -1.24 eV), charge transfer of to *CO, CO bond elongation, and change in OCO to 121-134, indicating CO activation. For a CuPt cluster, approximately equivalent to the experimental composition of 0.14Pt-0.6Cu, the CO adsorbs at both O vacancy and Pt sites. It is notable that CO strongly interacts with CuPt/TiO ( -1.2 eV) together with C-O bond elongation of 0.2 Å, OCO change to 121, and gained by *CO, thus promoting CO activation. The preference of *CO at Pt facilitates reaction with H* to lead CH formation pathway rather than CO desorption (=-1.78 eV). The optimal Pt-Cu composition leads to better activity than only Cu. For higher Pt concentration, the selectivity for CH was observed but with lower yield. The CO activation at the CuPt/TiO interface, and the stability of *CO and *CO at the interfacial Pt atoms can support the observed CH selectivity.
4 Conclusion
In summary, using unbiased MLIP+BHMC together with ab-initio DFT approaches, we demonstrate that CuPt/TiO systems strongly facilitate CO activation at the interfacial sites. The process is facilitated by charge transfer and direct interaction. The calculated CO adsorption energies are from -0.5 eV to -1.2 eV, depending upon Cu/Pt ratio. The VB-maximum frontier states spatially localized at interfacial atoms, suggesting these atoms would be active sites for chemical reactivity.
The interface plays a key role in influencing the chemical activity of CuPt/TiO systems. The *CO *CO+*O step, where oxygen fills the vacancy site and CO adsorbs at an interfacial Pt site, is energetically favorable. The other elementary steps can result in *CHOO, *CHO, *CHOH, *CHOH, *CH, and *CH intermediates. The interface, with optimal CuPt composition, bolsters the stability of *CO and *CO, suggesting higher activity and CH selectivity. Our simulations qualitatively agree with the experimental observation of photocatalytic conversion of CO to CH. The maximum CH yield of 0.3 mol for CuPt composition with 0.14 mol% Pt and 2.5 mol% Cu, can be correlated with CO activation and CH-intermediates formation. The photocatalytic performance degrades with the decrease of density of interfacial sites, suggesting the necessity of optimal surface area of interfacial regions for selective CO conversion to hydrocarbons.
Comprehensive studies of several intermediate structures together with energetics and kinetics of the possible reaction pathways could shed more light on reaction mechanisms. Nonetheless, exploring PES of main reactants (HO/CO) and several CH-containing intermediates with activation energy, vibrational entropy, zero-point energy are computationally intensive, and is beyond the scope of the current work. We hope our studies encourage pursuing further study on CuPt/P25 or similar systems and motivate for developing the sophisticated MLIP potential for simulations of these photocatalytic systems with the reactants and intermediates.
5 Methods
5.1 Ab-initio calculations
We performed the ab-initio DFT calculations using VASP [64] and QE [65] on AWS EC2 computing platform. The MLIP, as shall be discussed, was trained on VASP generated datasets and a few DFT calculations were performed and validated using QE simulations. We used the plane-wave basis set and pseudopotential approaches. For the exchange-correlation of electrons, we used the generalized-gradient approximation (GGA) in the form of Perdew–Burke-Ernzerhof (PBE) functional.[66] We used the projector-augmented wave (PAW) pseudopotential method [67] for describing electron-ion interactions. Here, we chose the rutile TiO(110) since it is the most stable among its various surfaces and extensively studied model system.[68] We constructed the 13-atom sub-nanometer-sized CuPt nanoclusters, denoted by CuPt. The motivation for the choice of magic-numbered clusters partly arises from several earlier computational studies, e.g.[69, 70, 71, 72, 73] and partly come from the computationally tractable combinatorial problem. Further details on DFT calculations can be found in supporting information (SI) section S3.1, on energetics calculations in section S2.2 and Bader charge analysis in section S2.3.
5.2 Machine learning interatomic potential
To solve a combinatorial problem relating to complexities of structures involving five different chemical elements, we built a machine learning model based on Allegro – a deep equivariant neural network architecture.[32] We use a 6 Å radial cutoff and 2 interaction blocks for the Allegro-based MLIP. A polynomial envelope with cutoff of and eight trainable Bessel functions are used for a basis expansion of radial distances. We restrict the maximum irreducible representation of SO(3) for internal rotational features (i.e., the maximum order of geometric tensor embeddings transforming like type-l spherical harmonics) to . For each , we allowed for both even and odd parity features. The two-body embeddings consist of 3 layers of dimensions (32, 64, 128). Within each interaction block, we use latent MLIPs consisting of 3 layers of dimensions (128, 128, 128). As nonlinear activations, sigmoid linear unit (SiLU) functions are used throughout the entire architecture. MLIP optimizers, learning rates and numerical precisions are discussed in the SI section S3.4.
We employ an iterative training approach following the creation of an initial potential. Structures generated from basin-hopping Monte Carlo (BHMC) simulations,[41, 44] accelerated by the trained MLIP, go through a workflow for identifying unique configurations (utilizing uniform manifold approximation and projection (UMAP) and K-means clustering). Single-point DFT calculations on configurations exhibiting high errors relative to MLIP predictions are incorporated into the training set. Iterations persist until convergence in terms of energy ( 1 meV/atom) and force errors (0.1 eV/Å) is achieved. Additional details about the method are provided in the SI (see section S3.5 and S3.6).
5.3 Sampling unbiased training
The MLIP training dataset, previously described, encompasses atom coordinates, energy, and forces of CO@CuPt/TiO systems. The (6x2)-TiO(110) substrate, featuring a single oxygen vacancy, interacts with CuPt nanoclusters ( to 13), and CO is positioned at various adsorption sites. Overall, the training data includes bulk atoms (TiO substrate) and surface atoms (consisting of 4,690 C and O atoms from CO and Pt/Cu atoms). As a result, the MLIP training inherently favors TiO atoms to reduce overall prediction errors for the predicted forces, sacrificing for surface Ti and O atoms as well as the atoms comprising the nanoparticles and the adsorbate, which are crucial for understanding the surface science and catalytic activity.
This bias can be quantified by defining a gaussian density function (GDF) for the encoding of an atomic environment that describes the radial and angular distribution (Symmetry functions G) of the neighboring atoms within a certain cutoff radius (see SI section S3.5). For an arbitrary G for an atomic species in the entire space, we define the Gaussian Density function ) as:
where is the Gaussian width and is the dimension of the symmetry function vector, is the total number of atoms in the entire dataset. The ranges between 0 and 1, where 0 means scarce training points whereas 1 represents abundant training points.
Addressing biases arising from the redundancy of atomic environments in the training set, we also acknowledge non-uniformity in training on atomic forces. The existing loss functions for MLIP training treat the absolute error in forces as constant, irrespective of force magnitude, leading to higher relative force errors for smaller values. To overcome the identified issues, we propose the use of an alternative weighted loss function in the MLIP model ensuring improved and uniform training for the ab-initio DFT data. The modified loss function takes the following form:
where B is a scaling factor used to tune the impact of low-force data on the loss function (chosen to be 10 here), determines the relative weight of force error with respect to the energy error in the loss function ( here), is a monotonically increasing function (modified sigmoid function), A is a normalizing constant (that makes the average of to be 1), and b and c are parameters that are fine-tuned for the balanced training (chosen to be 150 and 1.0, respectively).
5.4 Experimental details
5.4.1 Synthesis of CuPt nanoparticles decorated with P25
The CuPt nanoparticles decorated with P25 were synthesized using methods as reported by Sorcar et al.[13] Before synthesis of CuPt nanoparticles, the reduced P25 was initially prepared by mixing 200 mg of P25 with 30 mg of NaBH in a mortar and pestle. The powder was then heated in an inert atmosphere (Ar flow) at 350C for about an hour. Upon cooling, the P25 was purified by washing in water and ethanol and then the sample was centrifuged. This was repeated at least five times. The resulting powder was dried overnight in a vacuum oven at 100C.
The platinum was deposited onto the reduced P25 via photo-deposition. 40 mg of the P25 was initially dispersed in 10 mL of the 4:1 HO:CHOH solution. HPtCl solution was then added to the dispersion and allowed to stir for one hour. The dispersion was then irradiated by AM1.5 light by a 300 W Xe lamp at 1 sun intensity for two hours. The Pt-P25 was collected by centrifuging. Then it was washed with HO and ethanol mixture three times and dried under vacuum. Copper was subsequently deposited in a similar method using Cu(NO) instead of HPtCl. The CuPt-P25 samples were then deposited onto glass frits before loading into the photoreactor. Dispersions of the samples in isopropanol were drop-coated onto Aceglass glass frits with porosity C until about 15 mg was loaded. The frits were then dried under vacuum overnight.
5.4.2 Experimental setup for measuring photocatalytic activity of CuPt-P25
The catalyst coated glass frit was placed into a custom-built steel photoreactor fitted with a quartz window. Before testing, the reactor was purged by evacuating the reactor and refilling with 1:1 Ar:CO ten times. 1.5 sccm of Ar and 1.5 sccm CO were bubbled through a water filled gas washing bottle and fed into the reactor from the top, passing through the frit, and exiting the bottom of the reactor. The reactor was then irradiated by a 300 W Xe lamp at one sun intensity through an AM1.5 filter. The product gas composition was analyzed by a gas chromatograph (SRI GC MG5).